OpenStudy (thomas5267):

@Callisto Can you give me some more examples of doing integration "your way"?

OpenStudy (callisto):

\[\int tanx dx\]

OpenStudy (callisto):

\[=\int \frac{sinx}{cosx}dx\]\[=-\int \frac{1}{cosx}d(cosx)\]\[=-\ln|cosx|+C\] Try: \[\int cotx dx\]

OpenStudy (thomas5267):

\[ \begin{align*} &\int\cot(x)\,dx\\ =&\int\frac{1}{\tan(x)}\,dx\\ =&\int\frac{\cos(x)}{\sin(x)}\,dx\\ =&\int\frac{1}{\sin(x)}\,d(\sin(x))\\ =&\int\ln(\sin(x)) \end{align*} \] Seems legit...

OpenStudy (callisto):

No int for the last line \[=\int\frac{1}{sinx}d(sinx)\]\[=ln|sinx|+C\]